# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ 3.2.2 Polynomial Time Algorithm
In addition, the time and space complexity of the algorithm is at most polynomial in terms of the parameters $N,k,\epsilon,\frac{1}{\alpha}$.
@@
We must note here that our proof of Theorem 3.8 and the algorithm are inspired by the approach undertaken by Canonne et al., (2023). However, the algorithms and results of Canonne et al., (2023) only work in the region when $\left\lVert\mu\right\rVert_{2}\lesssim 1$ and there is no constraint under the alternative, i.e., $K=\mathbb{R}^{d}$. Therefore, we crucially modified the algorithm, including techniques to allow for $\mu$ of arbitrary (Euclidean) length and the use of the projection $P^{\star}$ to allow for the prior knowledge on $\mu$. We also fixed some flaws left in the argument of Canonne et al., (2023). We kindly refer readers to Appendix B.6 for the complete proof of Theorem 3.8.
@@ 3.3 Main Results
In this section, we integrate the preceding theorems and bounds and summarize the main results. The lower bound is synthesized from Theorems 3.2, 3.3, 3.4. Although the bounds are derived through different approaches, an interesting phase transition phenomenon emerges, determined by the ranges of $\epsilon$ over which each term dominates others.
@@
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{1}^{\star},k_{2}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 2.3 and 2.4.
In the proof of Corollary 3.12 below, we show that the first term, which is independent with $\epsilon$ [^5] We should recall that $k_{1}^{\star}$ is independent with $\epsilon$, dominates the other two if and only if $\epsilon\leq\frac{1}{\sqrt{N}}$. Similarly, the last term, which is independent with $N$, is the dominant if and only if $\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}$. Therefore, the set $\left\{\epsilon\left\lvert\right.0\leq\epsilon<c_{0},\epsilon\in\left[\frac{1}{\sqrt{N}},\sqrt{\frac{k_{2}^{\star}}{N}}\right]\right\}$ [^6] We express in this form because $k_{2}^{\star}$ also depends on $\epsilon$. serves as a threshold belt. When $\epsilon$ lies below this belt, it is too small for the adversary $\mathcal{C}$ to exert its influence, and the optimal minimax rate becomes essentially independent of $\epsilon$. In other words, it is free to contaminate the dataset in this region. In contrast, if $\epsilon$ lies above the belt (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), then the power of $\mathcal{C}$ can be fully utilized, to the extent that increasing the sample size $N$ does not improve the problem’s difficulty. When $\epsilon$ falls within the belt, the minimax rate is the result of the delicate balance between the adversary $\mathcal{C}$ and the testing procedures $\phi_{e},\phi_{p}$. Based on this argument, (3.6) can be equivalently written as
$\rho\gtrsim\left\{\begin{tabular}[]{lll}$\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$,&$\left(0\leq\epsilon\lesssim\frac{1}{\sqrt{N}}\right)$&(Free contamination phase),\\ $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$,&$\left(\frac{1}{\sqrt{N}}\lesssim\epsilon\lesssim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$&(Balance phase),\\ $\epsilon^{2}\sigma^{2}$,&$\left(\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$,&(Overly contaminated phase).\end{tabular}\right$. (3.7)
@@
For the upper bound, the theorem below is a natural downstream conclusion of Theorem 3.6 and 3.8.
@@
Given the same setting as in Theorem 3.10, if $\rho\gtrsim E(\epsilon,K,N,\\ \sigma^{2})$, there exists a test $\phi_{e}$ achieving type-I and type-II errors uniformly less than $\alpha$. Furthermore, if $\rho\gtrsim P(\epsilon,K,N,\sigma^{2})$ and the determination of $k_{1}^{\star}$ and $k_{2}^{\star}$ can be completed within polynomial time of $\left(N,d,\frac{1}{\epsilon}\right)$, there exists a test $\phi_{p}$ that is able to finish the testing process within polynomial time of $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$ while still achieving type-I and type-II errors uniformly less than $\alpha$. $E(\epsilon,K,N,\sigma^{2})$ and $P(\epsilon,K,N,\sigma^{2})$ are defined in (3.3) and (3.5) respectively.
@@
For Corollary 3.12, it involves some refined analysis with (3.6) and $E(\epsilon,K,N,\sigma^{2})$. We kindly refer readers to Appendix B.7 for the detailed arguments.
3.4 Extension to $\ell_{p}$ Norm Separation for $1\leq p<2$
@@
The classic results for the testing problem (1.2) can be naturally generalized to (3.8) when $1\leq p<2$ via the Hölder’s inequality (see Lemma A.17) by Ingster and Suslina, (2003). While Ingster’s approach cannot be directly transferred to our settings, we establish similar generalized results with the $p$-convex condition and accordingly adjusted optimal dimensions and projections. Note that in this section, we shall substitute the $\ell_{2}$ norm with the $\ell_{p}$ norm for the Kolmogorov $k$-width. Furthermore, to simplify the analysis, we consider the discretized version (see the remark after Definition 2.2).
@@
Definition 3.14 ($p$-convex set).
For $p\geq 1$, a set $K\subset\mathcal{X}$ is $p$-convex if
@@
$K^{p}:=\left\{\left(\left|\theta_{1}\right|^{p},\dots,\left|\theta_{d}\right|^{p}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\}.$
@@
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 3.16.
Undoubtedly, the lower bound (3.10) recovers (3.6) as expected when $p=2$. A significant difference from the $\ell_{2}$ norm separation is that we have another optimal dimension regarding the last term in (3.10). The intrinsic reason is that when covariance matrices are known and identical, the KL divergence between Gaussian distributions — an essential quantity we leverage in the proof of Theorem 3.4, depends on the $\ell_{2}$ norm between the means. When considering the $\ell_{p}$ norm separation, we require $k_{p,3}^{\star}$ to adapt to this discrepancy. We defer the proof of Theorem 3.17 to Appendix B.8.
In $\ell_{2}$ norm separation, the upper bound (3.3) matches the lower bound (3.6) except for logarithmic factors. We identify similar results in $\ell_{p}$ norm separation, which is summarized as the theorem below.
@@
If $\rho\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}$, then there exists a test to distinguish between $H_{0}$ and $H_{1}$ with uniformly small type-I and type-II errors. Moreover, the time complexity of the algorithm is polynomial degree in $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$. Note that in the expression above, we omit the polylogarithmic factors in $\left(N,k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star},\frac{1}{\alpha},\frac{1}{\epsilon}\right)$.
At first examination, Theorem 3.18 appears to be a direct and natural corollary of Theorem 3.11. However, a closer inspection reveals subtleties in its derivation. A crucial observation is that, though one can select the minimum $k_{p}^{\star}:=\min\limits\left\{k_{p,1}^{\star}k_{p,2}^{\star},k_{p,3}^{\star}\right\}$ and its optimal projection $P_{p}^{\star}$, and is guaranteed that there exist three vectors $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ such that $\left\lVert P_{p}^{\star}\theta_{p,1}\right\rVert_{2}^{2}\gtrsim\frac{\sqrt{k^{\star}}}{N}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,2}\right\rVert_{2}^{2}\gtrsim\epsilon\sqrt{\frac{k_{p}^{\star}}{N}}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,3}\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\sigma^{2}$, $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ are not necessary to be identical. In Theorem 3.6 or Theorem 3.8, however, we require a single vector satisfying all the conditions simultaneously. In fact, similar to the $\ell_{2}$ norm case, we do need to select the minimal optimal dimension. To overcome the difficulty specified above, we again need to exploit the properties of the $p$-convex orthosymmetric set and the Kolmogorov widths. The proof of Theorem 3.18 is provided in Appendix B.9.
@@ 4 Experiments
In this section, we present the results of simulation to verify the correctness of the theorems above and algorithms in Appendix C on the application side. We will generate artificial data, design different strategies for the adversary to contaminate the data, and run the (polynomial time) algorithm to check its effectiveness. The pseudo code is attached in Appendix C.
@@ 4.2 Choice of Constraints
where $k^{\star}=\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}$. We verify in Appendix B.10 that the approach (4.2) ensures that the resulting vector $\mu_{c}$ achieves the minimal value of $\left\lVert P^{\star}\mu\right\rVert_{2}$ among the set $\left\{v\left\lvert\right.v\in\mathbb{R}^{d},v\in K_{e},\left\lVert v\right\rVert_{2}=c\right\}$, where $P^{\star}$ is the optimal projection operator corresponding to $k^{\star}$. We believe that this property places $\mu_{c}$ in a near-worst-case position for the algorithm, and thus provides an appropriate setting to test its robustness.
@@ 11   Compute $\mu_{0}$ according to the constraint type
12  Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{arr}_{i}=\left\lVert\widetilde{X}_{i}-\mu_{0}\right\rVert_{2}$
@@ 4.4 Experimental Results
The experimental results are presented in the following figures. During the experiments, we ensured that the type-I error remained near zero, while identifying the empirical $\widetilde{\rho}_{\text{critical}}$ under $H_{1}$ as the critical value of $c$ when the corresponding type-II error exactly crosses $\alpha=0.05$.
Figure 3: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=200$.
Figure 4: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=2000$.
Figure 5: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=100$.
Figure 6: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=500$.
We observed several notable phenomena from the simulation results. First, the algorithm performs reasonably well across all parameter configurations: it successfully rejects $H_{0}$ when $\left\lVert\mu\right\rVert_{2}$ is sufficiently large and fails to reject when the signal strength is too weak, in both the classical and the high-dimensional regimes. Second, focusing on $\rho_{\text{critical}}$ — the critical norm of $\mu$ that yields a type-II error of $\alpha=0.05$ — we find that $\rho_{\text{critical}}$ decreases as $N$ increases, and increases as $d$ grows. This behavior is consistent with theoretical intuition: larger sample sizes improve detectability, whereas higher dimensionality makes the testing problem more challenging. However, a more refined observation is that the influence of $d$ diminishes once the dimension becomes sufficiently large (e.g., around $d=50$ in our experiments). The reason behind this phenomenon is that, although $d$ increases to relatively large values, the optimal projection dimension $k^{\star}$ becomes increasingly insensitive to $d$ under our setting of $N=1,\sigma=1$, and the ellipsoidal constraint. A direct calculation shows that when $N=2000$, the optimal dimensions corresponding to full dimensions $[5,10,20,50]$ are $[5,9,10,12]$, respectively. This demonstrates that the gaps between the optimal dimensions can be substantially smaller than those between the actual dimensions, which leads to the phenomenon. Finally, the plots indicate that as $N$ increases, the performance of the algorithm becomes less sensitive to variations in $\epsilon$. Though not fully verified, we believe that this is attributed to the fact that for larger $N$, the eigensubspace structure of $\mathbf{X}\mathbf{X}^{\top}$ (or $\mathbf{X}^{\top}\mathbf{X}$) becomes more stable with respect to the corrupted observations.

### v3

@@ 3.2.2 Polynomial-Time Algorithm
In addition, the time and space complexity of the algorithm is at most polynomial in terms of the parameters $N,k,\frac{1}{\epsilon},\frac{1}{\alpha}$.
@@
We acknowledge that the proof of Theorem 3.11 and the associated algorithm build upon the framework established in Canonne et al., (2023). However, the results in Canonne et al., (2023) are restricted to the regime where $\lVert\mu\rVert_{2}\lesssim 1$ in the unconstrained setting (i.e., $K=\mathbb{R}^{d}$). Consequently, we have significantly extended the algorithm to accommodate arbitrary signal magnitudes and to incorporate the prior constraint $K$ via the optimal projection $P^{\star}$. Furthermore, our analysis refines the arguments of Canonne et al., (2023) to address certain technical issues. We kindly refer readers to Appendix C.6 for the complete proof of the properties regarding the $\omega$-regularity and Theorem 3.11 and Appendix A for the pseudo code of the polynomial-time algorithm.
@@ 3.3 Main Results
In this section, we integrate the preceding theorems and bounds and summarize the main results. The lower bound is synthesized from Theorems 3.2, 3.5, 3.6. Although the bounds are derived through different approaches, an interesting phase transition phenomenon emerges, determined by the ranges of $\epsilon$ over which each term dominates others.
@@
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{1}^{\star},k_{2}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in Definition 2.3, 2.4.
Roughly speaking, Theorem 3.13 demonstrates that the minimax lower bound of the testing problem (1.2) is jointly determined by three parts — the geometric properties of the set $K$, which is represented by the factors $k_{1}^{\star},k_{2}^{\star}$; the properties of the original data, which is implied by $N$ and $\sigma$; and the corruption process, which is implied by $\epsilon$.
The corruption fraction $\epsilon$ connects the three components in the lower bound. In the proof of Corollary 3.15 below, we show that the first term, which is independent with $\epsilon$ [^5] We should recall that $k_{1}^{\star}$ is independent with $\epsilon$, dominates the other two if and only if $\epsilon\leq\frac{1}{\sqrt{N}}$. Similarly, the last term, which is independent with $N$, is the dominant if and only if $\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}$. Therefore, the set $\left\{\epsilon\left\lvert\right.0\leq\epsilon<c_{0},\epsilon\in\left[\frac{1}{\sqrt{N}},\sqrt{\frac{k_{2}^{\star}}{N}}\right]\right\}$ [^6] We express in this form because $k_{2}^{\star}$ also depends on $\epsilon$. serves as a threshold belt. When $\epsilon$ lies below this belt, it is too small for the adversary $\mathcal{C}$ to exert its influence, and the optimal minimax rate becomes essentially independent of $\epsilon$. In other words, it is free to contaminate the dataset in this region. In contrast, if $\epsilon$ lies above the belt (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), then the power of $\mathcal{C}$ can be fully utilized, to the extent that increasing the sample size $N$ does not improve the problem’s difficulty. When $\epsilon$ falls within the belt, the minimax rate is the result of the delicate balance between the adversary $\mathcal{C}$ and the testing procedures $\phi_{e},\phi_{p}$. Based on this argument, (3.8) can be equivalently written as
$\rho^{2}\gtrsim\left\{\begin{tabular}[]{lll}$\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$,&$\left(0\leq\epsilon\lesssim\frac{1}{\sqrt{N}}\right)$&(Free contamination phase),\\ $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$,&$\left(\frac{1}{\sqrt{N}}\lesssim\epsilon\lesssim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$&(Balance phase),\\ $\epsilon^{2}\sigma^{2}$,&$\left(\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$,&(Overly contaminated phase).\end{tabular}\right$. (3.9)
@@
For the upper bound, Theorem 3.14 below is a natural downstream conclusion of Theorem 3.8 and 3.11.
@@
Given the same setting as in Theorem 3.13, if $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$, there exists a test $\phi_{e}$ achieving type-I and type-II errors uniformly less than $\alpha$. Furthermore, if $\rho\gtrsim P(\epsilon,K,N,\sigma^{2})$ and the determination of $k_{1}^{\star}$ and $k_{2}^{\star}$ can be completed within polynomial-time of $\left(N,d,\frac{1}{\epsilon}\right)$, there exists a test $\phi_{p}$ that is able to finish the testing process within polynomial-time of $\left(N,k_{1}^{\star},k_{2}^{\star},\frac{1}{\epsilon},\frac{1}{\alpha}\right)$ while still achieving type-I and type-II errors uniformly less than $\alpha$. $E(\epsilon,K,N,\sigma^{2})$ and $P(\epsilon,K,N,\sigma^{2})$ are defined in (3.5) and (3.7) respectively.
@@
For Corollary 3.15, it involves some refined analysis with the main lower bound (3.8) and $E(\epsilon,K,N,\sigma^{2})$. We kindly refer readers to Appendix C.7 for the detailed arguments.
4 Extension to $\ell_{p}$ Norm Separation for $1\leq p<2$
@@
The classic results for the testing problem (1.2) can be naturally generalized to (4.1) when $1\leq p<2$ via the Hölder’s inequality by Ingster and Suslina, (2003). While Ingster’s approach cannot be directly transferred to our settings, we establish similar generalized results with the $p$-convex condition and accordingly adjusted optimal dimensions and projections. Note that in this section, we shall substitute the $\ell_{2}$ norm with the $\ell_{p}$ norm for the Kolmogorov $k$-width. Furthermore, to simplify the analysis, we only consider the discretized version (see the remark after Definition 2.2).
@@
Definition 4.2 ($p$-convex orthosymmetric (PCO) set).
For $p\geq 1$, a set $K\subset\mathcal{X}$ is a $p$-convex orthosymmetric set if
@@
$K^{p}:=\left\{\left(\left|\theta_{1}\right|^{p},\dots,\left|\theta_{d}\right|^{p}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\};$
(3), $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
It is not difficult to verify that the whole space, hyperrectangles and ellipsoids are still all PCO sets for $1\leq p<2$.
@@
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in Definition 4.4.
Undoubtedly, the lower bound (4.3) recovers (3.8) as expected when $p=2$. A significant difference from the $\ell_{2}$ norm separation is that we have another optimal dimension regarding the last term in (4.3). The intrinsic reason is that when covariance matrices are known and identical, the KL divergence between Gaussian distributions — an essential quantity we leverage in the proof of Theorem 3.6, depends on the $\ell_{2}$ norm between the means. When considering the $\ell_{p}$ norm separation, we require $k_{p,3}^{\star}$ to adapt to this discrepancy. We defer the proof of Theorem 4.5 to Appendix C.8.
In $\ell_{2}$ norm separation, the upper bound (3.5) matches the lower bound (3.8) except for logarithmic factors. We identify similar results in $\ell_{p}$ norm separation, which is summarized as Theorem 4.6 below.
@@
If $\rho\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}$, then there exists a test to distinguish between $H_{0}$ and $H_{1}$ with uniformly small type-I and type-II errors. Moreover, the time complexity of the algorithm is polynomial-time in $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$. Note that in the expression above, we omit the polylogarithmic factors in $\left(N,k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star},\frac{1}{\alpha},\frac{1}{\epsilon}\right)$.
At first examination, Theorem 4.6 appears to be a direct and natural corollary of Theorem 3.14. However, a closer inspection reveals subtleties in its derivation. A crucial observation is that, though one can select the minimum $k_{p}^{\star}:=\min\limits\left\{k_{p,1}^{\star}k_{p,2}^{\star},k_{p,3}^{\star}\right\}$ and its optimal projection $P_{p}^{\star}$, and is guaranteed that there exist three vectors $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ such that $\left\lVert P_{p}^{\star}\theta_{p,1}\right\rVert_{2}^{2}\gtrsim\frac{\sqrt{k^{\star}}}{N}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,2}\right\rVert_{2}^{2}\gtrsim\epsilon\sqrt{\frac{k_{p}^{\star}}{N}}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,3}\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\sigma^{2}$, $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ are not necessary to be identical. In Theorem 3.8 or Theorem 3.11, however, we require a single vector satisfying all the conditions simultaneously. In fact, similar to the $\ell_{2}$ norm case, we do need to select the minimal optimal dimension. To overcome the difficulty specified above, we again need to exploit the properties of the $p$-convex orthosymmetric set and the Kolmogorov widths. The proof of Theorem 4.6 is provided in Appendix C.9.
@@ 5 Experiments
In this section, we present the results of simulation to verify the correctness of the theorems above and algorithms in Appendix A on the application side. We generate artificial data, design different strategies for the adversary to contaminate the data, and run the (polynomial-time) algorithm to check its effectiveness. The pseudo code is attached in Appendix A.
@@ 5.2 Choice of Constraints
where $k^{\star}=\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}$. We verify in Appendix C.10 that the approach (5.2) ensures that the resulting vector $\mu_{c}$ achieves the minimal value of $\left\lVert P^{\star}\mu\right\rVert_{2}$ among the set $\left\{v\left\lvert\right.v\in\mathbb{R}^{d},v\in K_{e},\left\lVert v\right\rVert_{2}=c\right\}$, where $P^{\star}$ is the optimal projection operator corresponding to $k^{\star}$. We believe that this property places $\mu_{c}$ in a near-worst-case position for the algorithm, and thus provides an appropriate setting to test its robustness.
@@ 15 $\widetilde{\mathbf{X}}_{C}=\mathbf{X}_{C}$
6 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{dist}$
@@ 11   Compute $\mu_{0}$ according to the constraint type
12  Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{dist}_{i}=\left\lVert\widetilde{X}_{i}-\mu_{0}\right\rVert_{2}$
13 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{dist}$
@@ 5.4 Experimental Results
Experimental results for $N=200$ in the classical case and $N=100$ in the high-dimensional case are presented in the following figures. Results for other parameter configurations are similar and are deferred to Appendix A. In our experiments, we maintained the type-I error near zero while identifying the empirical $\widetilde{\rho}_{\text{critical}}$ under $H_{1}$ as the critical value of $\rho$ when the corresponding type-II error exactly crosses $\beta=0.05$.
Figure 3: Top: Rejection rate (or Empirical power) versus the $\ell_{2}$ norm of the mean under $H_{1}$ in the classic setting ($N=200$). Bottom: Rejection rate versus the $\ell_{2}$ norm of the mean in the high-dimensional setting ($N=100$).
We observed several notable phenomena from the simulation results. First, the algorithm performs reasonably well across all parameter configurations, even for the moderate sample sizes: it successfully rejects $H_{0}$ when $\left\lVert\mu\right\rVert_{2}$ is sufficiently large and fails to reject when the signal strength is too weak, in both the classical and the high-dimensional regimes. Second, focusing on $\rho_{\text{critical}}$ — the critical norm of $\mu$ that yields a type-II error of $\alpha=0.05$ — we find that $\rho_{\text{critical}}$ decreases as $N$ increases, and increases as $d$ grows. This behavior is consistent with theoretical intuition: larger sample sizes improve detectability, whereas higher dimensionality makes the testing problem more challenging. However, a more refined observation is that the influence of $d$ diminishes once the dimension becomes sufficiently large (e.g., around $d=50$ in our experiments). The reason behind this phenomenon is that, although $d$ increases to relatively large values, the optimal projection dimension $k^{\star}$ becomes increasingly insensitive to $d$ under our setting of $N=1,\sigma=1$, and the ellipsoidal constraint. A direct calculation shows that when $N=2000$, the optimal dimensions corresponding to full dimensions $[5,10,20,50]$ are $[5,9,10,12]$, respectively. This demonstrates that the gaps between the optimal dimensions can be substantially smaller than those between the actual dimensions, which leads to the phenomenon. Finally, the plots indicate that as $N$ increases, the performance of the algorithm becomes less sensitive to variations in $\epsilon$. Though not fully verified, we believe that this is attributed to the fact that for larger $N$, the eigensubspace structure of $\mathbf{X}\mathbf{X}^{\top}$ (or $\mathbf{X}^{\top}\mathbf{X}$) becomes more stable with respect to the corrupted observations.